We finished the entrance to drill and science trenches on a cold and windy morning.
The white weatherport on top of the stairs and elevator to the drill and science trench.
This morning began cold and windy. With temperatures from -30°C to -25°C and 15 knot wind in the morning, our newcomers got an experience of how weather normally is in the beginning of May. A valiant crew managed to erect a white weatherport on top of the elevator and staircase. The wind made it complicated to put the cover on; but as soon as the cover was on, the wind dropped and temperatures rose to make the afternoon really pleasant. Weather up here is not without a certain irony.
At 16.00 the weather pulled another trick out of its sleeve: A wall of cloud came rolling in and engulfed the camp in cloud and snow, much too similar to the weather of last Saturday to Thursday for our taste. This time however cloud cover only lasted two hours, so at 18.00 the sky was clear again. With tent cover on stair case and elevator, we will soon close the inclined trench that leads down to the bottom of the stair case. Peter worked hard on covering the trench with roof. Friday night, he worked until midnight. He is a very hard worker. The roof covered trench will serve as storage for frozen food – "the cooks freezer".
What we have done today:
- Drilling with the NEEM long drill.
- Logging. Last bag: 477. Depth: 262.35 m.
- Erecting white Weather Port over elevator and stair case. A perfect fit.
- Cleaning up in camp.
- Extending roof of inclined trench.
- Construction of shelves in inclined trench for cooks freezer.
- Repaired a harddisk error in the line scanner computer.
- Saturday night. Henriette, Anna W. and Celia were Saturday night cooks, and we had a nicely spiced meat stew with rice and Portobello mushrooms. For dessert: Roasted apples with nuts and caramel sauce.
Ad.1: This is how the drillers report looks today: An interesting day in the drill trench. The ring above the pump was removed in order to test the suspicion that it could block the finest chips from leaving the pump. The next run gave a 3 m core with constant current until at the very end. ….high current due to packing of the pump. The next run was quite stable until the lower half of the chips chamber was packed, resulting in 2.8 m long core. We had often observed that the mid shaft polyethylene type booster could be quite inefficient in moving chips to the upper half of the chips chamber. Thus we replaced it with a shiny and smooth 1 turn brass booster. The next run, 3 m long, gave stable current until the end, no packed pump. Pull out was quite easy. Perhaps we are tracking down a stable mode of drilling. The pitch is, however, amazingly low, 2.5 to 3.9 mm, producing very fine chips, which the drill seems to accept after the changes made during the day. Chips recovery is 100 % and inclination unchanged, close to 1 deg. The last run was hampered by loose ice chunks in the hole. We drilled 13.15 m core in 5 runs; drillers depth 249.62 m.
Ad.7: Jeppe has just made it to camp, when he had to repair a severe harddisk problem on the line scanner. He fixed it, and now the line scanner computer is running.
Weather: Fine most of the day, with a brief, passing but significant change. -30°C to -14°C, 5-15 knots mainly from ESE. Visibility: Unrestricted. Weather was clear until 16.00 as a wall of clouds rolled in. It suddenly was over cast with snow, and the wind turned from ESE to SW. At 18.00 weather cleared again, and the wind turned back to ESE.
Field leader, J.P. Steffensen
|← Previous entry||Next entry →|